# Illuminating the Physics of Dark Energy with the Discovery Simulations

### v1

@@ 1 Introduction
With the arrival of data from the Dark Energy Spectroscopic Instrument (DESI, DESI Collaboration et al., 2016, 2022), we are now in the era of Stage IV dark energy experiments. The analysis of the Baryonic Acoustic Oscillation (BAO) signature measured from the first year of DESI observations (DESI Y1, DESI Collaboration et al., 2024a) suggests that the cosmic acceleration might not be driven by a simple cosmological constant, $\Lambda,$ but instead by a cosmic fluid with an equation of state parameter, $w$, that evolves in time. If confirmed, this finding would have profound implications for cosmology, as these two scenarios for accounting for the cosmic acceleration have dramatically different physical interpretations. However, the initial results from DESI Y1 are inconclusive on this question. Future analyses of DESI data will improve both statistical precision and treatment of systematic errors, and so it is entirely possible that BAO measurements of DESI Y3 will favor $\Lambda$CDM over evolving dark energy; it is also possible that the BAO signal in DESI is simply unable to conclusively discriminate between these two scenarios, and that alternative cosmological probes are needed to shed further light on the physics of dark energy.
@@
Progress in the quality of cosmological N-body simulations has been especially rapid over the last decade. The simplest marker of this progress is the mass resolution that characterizes the simulations used by Stage IV relative to Stage III experiments. Within DESI, the AbacusSummit simulations (Maksimova et al., 2021) span a wide range of cosmological parameters, and the primary suite of boxes has $L_{\rm box}\approx 2$$h^{-1}\mathrm{Gpc}$ and $m_{\rm ptcl}\approx 2\times 10^{9}$$h^{-1}\mathrm{M}_{\odot}$. Within the Vera C. Rubin Observatory Legacy Survey of Space and Time (LSST [^1] https://rubinobservatory.org, Ivezić et al., 2019), the N-body simulation Outer Rim (Heitmann et al., 2019) with $L_{\rm box}=3$$h^{-1}\mathrm{Gpc}$ and $m_{\rm ptcl}=1.85\times 10^{9}$ $h^{-1}\mathrm{M}_{\odot}$ was used in the production of the DC2 synthetic galaxy catalog (LSST Dark Energy Science Collaboration et al.(2021)LSST Dark Energy Science Collaboration (LSST DESC), Abolfathi, Alonso, Armstrong, Aubourg, Awan, Babuji, Bauer, Bean, Beckett, Biswas, Bogart, Boutigny, Chard, Chiang, Claver, Cohen-Tanugi, Combet, Connolly, Daniel, Digel, Drlica-Wagner, Dubois, Gangler, Gawiser, Glanzman, Gris, Habib, Hearin, Heitmann, Hernandez, Hložek, Hollowed, Ishak, Ivezić, Jarvis, Jha, Kahn, Kalmbach, Kelly, Kovacs, Korytov, Krughoff, Lage, Lanusse, Larsen, Le Guillou, Li, Longley, Lupton, Mandelbaum, Mao, Marshall, Meyers, Moniez, Morrison, Nomerotski, O’Connor, Park, Park, Peloton, Perrefort, Perry, Plaszczynski, Pope, Rasmussen, Reil, Roodman, Rykoff, Sánchez, Schmidt, Scolnic, Stubbs, Tyson, Uram, Villarreal, Walter, Wiesner, Wood-Vasey, & Zuntz, LSST DESC). The Outer Rim was also used by the OpenUniverse collaboration to produce the Roman-Rubin image simulation (OpenUniverse et al., 2025) in collaborative work between LSST and the Nancy Grace Roman Space Telescope [^2] https://roman.gsfc.nasa.gov. Within the Euclid collaboration [^3] https://www.cosmos.esa.int/web/euclid (Laureijs et al., 2011), the flagship simulation (Euclid Collaboration et al., 2024) has $L_{\rm box}=3.6$$h^{-1}\mathrm{Gpc}$ and $m_{\rm ptcl}=1\times 10^{9}$$h^{-1}\mathrm{M}_{\odot}$.
@@
Motivated by these developments, in this paper we present the Discovery simulations: two high-resolution, Gpc-scale cosmological N-body simulations that we have designed to provide a testbed for studies of dark energy that follow up on recent DESI results. The Discovery simulations are a pair of boxes with identical initial conditions; the only difference between the two simulations is cosmology. The first box uses a $\Lambda$CDM cosmology, while the second box contains a dark energy equation of state $w$ which evolves in time. The particular values of the cosmological parameters in each simulation were chosen to closely mimic the best-fitting cosmologies in the DESI Y1 BAO analyses (see §2 for details). Data products for the Discovery simulations include halo catalogs as well as substructure merger trees.
@@ 2 Simulations
Figure 2: Visual comparison of a small region in the simulations at $z=0$. Left: $\Lambda\mathrm{CDM}$; Right: $w_{0}w_{a}\mathrm{CDM}$. The differences are subtle but still clearly visible at the substructure level. This comparison demonstrates the precision cosmology challenge to obtain cosmological measurements that are sensitive to the smallest changes in structure formation. A short movie showing the evolution of a small subvolume is available here.
@@ 2.1 Parameter Choices
Each simulation covers a volume of (1500Mpc)3 and evolves 67203 particles with a force resolution of $\sim 2$kpc. We begin the simulations at $z=200$ and use the Zel’dovich approximation to assign initial particle positions and velocities (Zel’dovich, 1970). We generate input transfer functions with CAMB (Lewis et al., 2000). We assume a flat cosmology and massless neutrinos, $\omega_{\nu}=0$.
@@
By combining the large, survey-scale volumes of the simulations with high-resolution in mass, we are able to extract substructure merger trees and build detailed synthetic skies. These simulations therefore provide a unique resource for the cosmology community to further investigate measurements obtained from ongoing and upcoming surveys. In Fig. 2 we compare a small region of the two simulations and illustrate the subtle differences in substructure due to the small differences in cosmological parameter choices. This image presents a visual glimpse of the challenge for precision cosmology, and demonstrates the importance of high resolution, without which all differences in structure would be washed out.
@@ 2.2 Merger Trees and Substructure Tracking
In this work, we use the core merger trees from the two Discovery simulations to investigate the impact of an evolving dark energy on the mass assembly history of halos (see § 3.3), and subsequently how differences in halo mass assembly history manifest themselves as differences in the galaxy population (see § 3.4).
@@ 3.1 Matter Power Spectra
The nonlinear matter power spectrum and its evolution over time hold valuable cosmological information. HACC allows for a fast evaluation of the power spectrum on-the-fly, and as part of the simulation runs we have evaluated the power spectrum at many time steps. The power spectrum is not directly measured from the particles, but rather by first generating a density field on a grid using a cloud-in-cell assignment and applying a discrete FFT to obtain $\delta({\bf k})$. From this we determine $P({\bf k})=|\delta({\bf k})|^{2}$ which is then binned in amplitudes to obtain the matter power spectrum $P(k)$.
@@ 3.2 Halo Mass Functions
Figure 5: Influence of $w_{0}w_{a}\mathrm{CDM}$ cosmology on halo accretion rate. The fractional difference between the halo mass accretion rate in the $w_{0}w_{a}\mathrm{CDM}$ simulation compared to the $\Lambda\mathrm{CDM}$ simulation is plotted as a function of redshift in several halo mass bins between $M_{h}(z=0)=11$ $\mathrm{M}_{\odot}$ (blue) and $M_{h}(z=0)=14.5$ $\mathrm{M}_{\odot}$ (red).
@@ 3.3 Mass Assembly of Halos
In Fig. 5 we show the fractional difference between the halo mass accretion rate in the $w_{0}w_{a}\mathrm{CDM}$ and $\Lambda\mathrm{CDM}$ simulations as a function of redshift in several halo mass bins between $M_{h}(z=0)=11$ $\mathrm{M}_{\odot}$ (blue) and $M_{h}(z=0)=14.5$ $\mathrm{M}_{\odot}$ (red). At low redshift ($z<0.5$), $w_{0}w_{a}\mathrm{CDM}$ halos exhibit a slightly faster mass accretion rate than $\Lambda\mathrm{CDM}$ halos, with $10^{11}$ $\mathrm{M}_{\odot}$ halos experiencing a $\sim$4higher accretion rate at $z=0$ in the $w_{0}w_{a}\mathrm{CDM}$ simulation. Higher mass halos, while still experiencing a faster accretion rate at $z=0$, exhibit smaller deviations from $\Lambda\mathrm{CDM}$ (1-2%). These deviations from $\Lambda\mathrm{CDM}$ at low redshift are to be expected given the advent of the dark energy dominated regime at $z<1$. At redshifts greater than $z=1$ the difference between the halo mass accretion rates of low mass halos in the two simulations is very small.
@@ 3.4 Star Formation Histories of Galaxies
Figure 6: Influence of $w_{0}w_{a}\mathrm{CDM}$ cosmology on star formation rate. The fractional difference between the star formation rate in the $w_{0}w_{a}\mathrm{CDM}$ simulation compared to the $\Lambda\mathrm{CDM}$ simulation is plotted as a function of redshift in several halo mass bins between $M_{h}(z=0)=11$ $\mathrm{M}_{\odot}$ (blue) and $M_{h}(z=0)=14.5$ $\mathrm{M}_{\odot}$ (red).
We use this novel galaxy-halo connection model to populate each of the Discovery simulations with galaxies. In Fig. 6 we show the fractional difference in star formation rate for central galaxies in the $w_{0}w_{a}\mathrm{CDM}$ and $\Lambda\mathrm{CDM}$ simulations as a function of redshift in several halo mass bins between $M_{h}(z=0)=11$ $\mathrm{M}_{\odot}$ (blue) and $M_{h}(z=0)=14.5$ $\mathrm{M}_{\odot}$ (red). In this figure, the same Diffstarpop model is applied to both simulations; in other words, differences in the star formation rates are due to differences in cosmology and halo mass assembly, and not due to differences in star formation parameters.
@@
The trends shown in Fig. 6 are driven by a combination of three factors: the differences in baryon fraction between the two cosmologies, the differences in halo mass assembly rates between the two cosmologies, and the particular star formation model that was applied to the Discovery simulations in this work. The Diffstar model assumes that $\mathrm{d}M_{\mathrm{g}}/\mathrm{d}t\propto f_{\mathrm{b}}\times\mathrm{d}M_{\mathrm{h}}/\mathrm{d}t$, where $\mathrm{d}M_{\mathrm{g}}/\mathrm{d}t$ is the accretion rate of baryonic material that is available for star formation, and $\mathrm{d}M_{\mathrm{h}}/\mathrm{d}t$ is the growth rate of total halo mass. The baryon fraction in the Discovery simulations is 0.1587 in the $\Lambda\mathrm{CDM}$ simulation and 0.1558 in the $w_{0}w_{a}\mathrm{CDM}$ simulations, a difference of $\sim 2$%. This difference in baryon fraction explains the overall $\sim 2$% offset in star formation rates between the two simulations at $z>1$, with the $w_{0}w_{a}\mathrm{CDM}$ cosmology exhibiting a slower star formation rate due to its lower baryon fraction. Had the Discovery simulations been run with the same baryon fraction, we would likely see no substantial offset in star formation rates at high redshift, and discrepancies at lower redshift would become more apparent.
@@ Acknowledgements
We thank Salman Habib, Eve Kovacs, Patricia Larsen, Matthew Becker, and other members of the Argonne CPAC group, as well as Johannes Lange, for useful discussions related to this work. We also thank Silvio Rizzi for visualizing the Discovery simulations. Work done at Argonne was supported under the DOE contract DE-AC02-06CH11357. We gratefully acknowledge use of the Bebop, Improv, and Swing supercomputers in the Laboratory Computing Resource Center at Argonne National Laboratory. The Discovery simulations were generated using resources of the Argonne Leadership Computing Facility, a U.S. Department of Energy (DOE) Office of Science user facility at Argonne National Laboratory and is based on research supported by the U.S. DOE Office of Science-Advanced Scientific Computing Research Program. This work was partially supported by the OpenUniverse effort, which is funded by NASA under JPL Contract Task 70-711320, ‘Maximizing Science Exploitation of Simulated Cosmological Survey Data Across Surveys’. The project that gave rise to these results received the support of a fellowship from ”la Caixa” Foundation (ID 100010434). The fellowship code is LCF/BQ/PI23/11970028.

### v2

@@ 1 Introduction
With the arrival of data from the Dark Energy Spectroscopic Instrument (DESI, DESI Collaboration et al., 2016, 2022), we are now in the era of Stage IV dark energy experiments. The analysis of the Baryonic Acoustic Oscillation (BAO) signature measured from the first year of DESI observations (DESI Y1, DESI Collaboration et al., 2024a) suggests that the cosmic acceleration might not be driven by a simple cosmological constant, $\Lambda,$ but instead by a cosmic fluid with an equation of state parameter, $w$, that evolves in time. This preference for an evolving dark energy model, $w_{0}w_{a}\mathrm{CDM}$, has also been found by the Dark Energy Survey (DES, DES Collaboration et al., 2025). If confirmed, these findings would have profound implications for cosmology, as these two scenarios for accounting for the cosmic acceleration have dramatically different physical interpretations. However, the initial results from DESI Y1 are inconclusive on this question. Future analyses of DESI data will improve both statistical precision and treatment of systematic errors, and so it is entirely possible that BAO measurements of DESI Y3 will favor $\Lambda$CDM over evolving dark energy; it is also possible that the BAO signal in DESI is simply unable to conclusively discriminate between these two scenarios, and that alternative cosmological probes are needed to shed further light on the physics of dark energy.
@@
Progress in the quality of cosmological N-body simulations has been especially rapid over the last decade. The simplest marker of this progress is the mass resolution that characterizes the simulations used by Stage IV relative to Stage III experiments. Within DESI, the AbacusSummit simulations (Maksimova et al., 2021) span a wide range of cosmological parameters, and the primary suite of boxes has a box size of $L_{\rm box}\approx 2$$h^{-1}\mathrm{Gpc}$ and a particle mass of $m_{\rm ptcl}\approx 2\times 10^{9}$$h^{-1}\mathrm{M}_{\odot}$. Within the Vera C. Rubin Observatory Legacy Survey of Space and Time (LSST [^1] https://rubinobservatory.org, Ivezić et al., 2019), the N-body simulation Outer Rim (Heitmann et al., 2019) with $L_{\rm box}=3$$h^{-1}\mathrm{Gpc}$ and $m_{\rm ptcl}=1.85\times 10^{9}$$h^{-1}\mathrm{M}_{\odot}$ was used in the production of the DC2 synthetic galaxy catalog (LSST Dark Energy Science Collaboration et al.(2021)LSST Dark Energy Science Collaboration (LSST DESC), Abolfathi, Alonso, Armstrong, Aubourg, Awan, Babuji, Bauer, Bean, Beckett, Biswas, Bogart, Boutigny, Chard, Chiang, Claver, Cohen-Tanugi, Combet, Connolly, Daniel, Digel, Drlica-Wagner, Dubois, Gangler, Gawiser, Glanzman, Gris, Habib, Hearin, Heitmann, Hernandez, Hložek, Hollowed, Ishak, Ivezić, Jarvis, Jha, Kahn, Kalmbach, Kelly, Kovacs, Korytov, Krughoff, Lage, Lanusse, Larsen, Le Guillou, Li, Longley, Lupton, Mandelbaum, Mao, Marshall, Meyers, Moniez, Morrison, Nomerotski, O’Connor, Park, Park, Peloton, Perrefort, Perry, Plaszczynski, Pope, Rasmussen, Reil, Roodman, Rykoff, Sánchez, Schmidt, Scolnic, Stubbs, Tyson, Uram, Villarreal, Walter, Wiesner, Wood-Vasey, & Zuntz, LSST DESC). The Outer Rim was also used by the OpenUniverse collaboration to produce the Roman-Rubin image simulation (OpenUniverse et al., 2025) in collaborative work between LSST and the Nancy Grace Roman Space Telescope [^2] https://roman.gsfc.nasa.gov. Within the Euclid collaboration [^3] https://www.cosmos.esa.int/web/euclid (Laureijs et al., 2011), the flagship simulation (Euclid Collaboration et al., 2024) has $L_{\rm box}=3.6$$h^{-1}\mathrm{Gpc}$ and $m_{\rm ptcl}=1\times 10^{9}$$h^{-1}\mathrm{M}_{\odot}$.
@@
Hydrodynamical simulations provide another method for modeling the formation and evolution of galaxies. These simulations include parameters that capture baryonic physics via phenomenological subgrid models. However, hydrodynamical codes with different implementations of subgrid models and choices of modeling parameters can exhibit substantial disagreement in results. Additionally, the choices of modeling parameters must be adjusted each time the resolution of the simulation is changed, in order to continue to obtain accurate agreement with observations used for calibration. The high computational cost of these simulations presents an additional difficulty, limiting our ability to use them for cosmological inference. For these reasons, the generation of synthetic sky catalogs for cosmological surveys still relies mostly on gravity-only simulations and flexible modeling of the galaxy-halo connection in post-processing.
Motivated by the recent DESI results, as well as recent developments in cosmological simulations, in this paper we present the Discovery simulations: two high-resolution, Gpc-scale cosmological N-body simulations that we have designed to provide a testbed for studies of dark energy that follow up on recent DESI results. The Discovery simulations are a pair of boxes with identical initial conditions; the only difference between the two simulations is cosmology. The first box uses a $\Lambda$CDM cosmology, while the second box contains a dark energy equation of state $w$ which evolves in time. The particular values of the cosmological parameters in each simulation were chosen to closely mimic the best-fitting cosmologies in the DESI Y1 BAO analyses (see §2 for details). Data products for the Discovery simulations include halo catalogs as well as substructure merger trees.
@@ 2.1 Parameter Choices
In this section we briefly discuss our choices for the simulation settings and cosmological parameters of the Discovery Simulations. In contrast to hydrodynamical simulations, gravity-only simulations have no free modeling parameters. Settings for, e.g., resolution, simulation volume, and starting redshifts, have been extensively discussed in the literature to achieve sub-percent level accuracy to at least $k\sim 1$Mpc-1 and high accuracy beyond (see, e.g., Heitmann et al., 2010; Schneider et al., 2016; Euclid Collaboration et al., 2021). Additionally, code comparisons have been performed in multiple studies (e.g., Heitmann et al., 2008; Grove et al., 2022). A recent summary of these comparisons can be found in Angulo & Hahn (2022).
The simulation parameters used in the Discovery Simulations have been carefully chosen to meet the accuracy requirements derived in these recent publications. Each simulation covers a volume of (1500Mpc)3 and evolves 67203 particles with a force resolution of $\sim 2$kpc. We begin the simulations at $z=200$ and use the Zel’dovich approximation to assign initial particle positions and velocities (Zel’dovich, 1970). We generate input transfer functions with CAMB (Lewis et al., 2000). Our approach for generating initial conditions is described in Heitmann et al. (2010), which found that the Zel’dovich approximation and the 2LPT approach yield very similar initial displacements and velocities at high redshift, and converge at $z=0$.
@@
By combining the large, survey-scale volumes of the simulations with high-resolution in mass, we are able to extract substructure merger trees and build detailed synthetic skies. These simulations therefore provide a unique resource for the cosmology community to further investigate measurements obtained from ongoing and upcoming surveys. In Fig. 2 we compare a small region of the two Discovery simulations to illustrate the subtle differences in substructure due to the small differences in cosmological parameter choices. This image presents a visual glimpse of the challenge for precision cosmology, and demonstrates the importance of high resolution simulations for studying differences in structure formation on small scales.
Figure 2: A visual comparison of a small region in the simulations at $z=0$. Left: $\Lambda\mathrm{CDM}$ (blue); Center: $w_{0}w_{a}\mathrm{CDM}$ (yellow); Right: an overlay of the two simulations. The differences are subtle but still clearly visible at the substructure level, particularly when overlaid. This comparison demonstrates the precision cosmology challenge to obtain cosmological measurements that are sensitive to the smallest changes in structure formation. A short movie showing the evolution of a small subvolume of the simulation is available here.
@@ 2.2 Merger Trees and Substructure Tracking
In this work, we use the core merger trees from the two Discovery simulations to investigate the impact of an evolving dark energy on the mass assembly history of halos (see § 3.3), and subsequently how differences in halo mass assembly history manifest themselves as differences in the galaxy population, using a novel model of the galaxy-halo connection (see § 3.4).
@@ 3.1 Matter Power Spectra
The nonlinear matter power spectrum and its evolution over time hold valuable cosmological information. HACC allows for a fast evaluation of the power spectrum on-the-fly, and as part of the simulation runs we have evaluated the power spectrum at many time steps. The power spectrum is not directly measured from the particles, but rather by first generating a density field on a grid using a cloud-in-cell assignment (appropriately normalized by the volume of the simulation) and applying a discrete FFT to obtain $\delta({\bf k})$. From this we determine the three-dimensional $P({\bf k})=|\delta({\bf k})|^{2}$ which is then binned in amplitudes and divided by the mode multiplicity in each bin to obtain the one-dimensional matter power spectrum $P(k)$. The smoothing from the CIC assignment is compensated for by dividing the power spectrum by the window function that corresponds to the CIC assignment scheme.
@@ 3.2 Halo Mass Functions
Figure 5: Influence of $w_{0}w_{a}\mathrm{CDM}$ cosmology on halo accretion rate. The fractional difference between the average halo mass accretion rate in the $w_{0}w_{a}\mathrm{CDM}$ simulation compared to the $\Lambda\mathrm{CDM}$ simulation is plotted as a function of redshift in several halo mass bins between $M_{h}(z=0)=11$ $\mathrm{M}_{\odot}$ (blue) and $M_{h}(z=0)=14.5$ $\mathrm{M}_{\odot}$ (red).
@@ 3.3 Mass Assembly of Halos
In Fig. 5 we show the fractional difference between the average halo mass accretion rate in the $w_{0}w_{a}\mathrm{CDM}$ and $\Lambda\mathrm{CDM}$ simulations as a function of redshift in several halo mass bins between $M_{h}(z=0)=11$ $\mathrm{M}_{\odot}$ (blue) and $M_{h}(z=0)=14.5$ $\mathrm{M}_{\odot}$ (red). At low redshift ($z<0.5$), $w_{0}w_{a}\mathrm{CDM}$ halos exhibit a slightly faster mean mass accretion rate than $\Lambda\mathrm{CDM}$ halos, with $10^{11}$ $\mathrm{M}_{\odot}$ halos experiencing a $\sim$4higher accretion rate at $z=0$ in the $w_{0}w_{a}\mathrm{CDM}$ simulation. Higher mass halos, while still experiencing a faster accretion rate at $z=0$, exhibit smaller deviations from $\Lambda\mathrm{CDM}$ (1-2%). These deviations from $\Lambda\mathrm{CDM}$ at low redshift are to be expected given the advent of the dark energy dominated regime at $z<1$. At redshifts greater than $z=1$ the difference between the mean halo mass accretion rates of low mass halos in the two simulations is very small.
@@ 3.4 Star Formation Histories of Galaxies
Figure 6: Influence of $w_{0}w_{a}\mathrm{CDM}$ cosmology on star formation rate. The fractional difference between the average star formation rate in the $w_{0}w_{a}\mathrm{CDM}$ simulation compared to the $\Lambda\mathrm{CDM}$ simulation is plotted as a function of redshift in several halo mass bins between $M_{h}(z=0)=11$ $\mathrm{M}_{\odot}$ (blue) and $M_{h}(z=0)=14.5$ $\mathrm{M}_{\odot}$ (red).
We use this novel galaxy-halo connection model to populate each of the Discovery simulations with galaxies. In Fig. 6 we show the fractional difference in average star formation rate for central galaxies in the $w_{0}w_{a}\mathrm{CDM}$ and $\Lambda\mathrm{CDM}$ simulations as a function of redshift in several halo mass bins between $M_{h}(z=0)=11$ $\mathrm{M}_{\odot}$ (blue) and $M_{h}(z=0)=14.5$ $\mathrm{M}_{\odot}$ (red). In this figure, the same Diffstarpop model is applied to both simulations; in other words, differences in the mean star formation rates are due to differences in cosmology and halo mass assembly, and not due to differences in star formation parameters.
@@
The trends shown in Fig. 6 are driven by a combination of three factors: the differences in baryon fraction between the two cosmologies, the differences in halo mass assembly rates between the two cosmologies, and the particular star formation model that was applied to the Discovery simulations in this work. The Diffstar model assumes that $\mathrm{d}M_{\mathrm{g}}/\mathrm{d}t\propto f_{\mathrm{b}}\times\mathrm{d}M_{\mathrm{h}}/\mathrm{d}t$, where $\mathrm{d}M_{\mathrm{g}}/\mathrm{d}t$ is the accretion rate of baryonic material that is available for star formation, $\mathrm{d}M_{\mathrm{h}}/\mathrm{d}t$ is the growth rate of total halo mass, and $f_{\mathrm{b}}$ is the baryon fraction in the simulation. The baryon fraction in the Discovery simulations is 0.1587 in the $\Lambda\mathrm{CDM}$ cosmology and 0.1558 in the $w_{0}w_{a}\mathrm{CDM}$ cosmology, a difference of $\sim 2$%. This difference in baryon fraction explains the overall $\sim 2$% offset in star formation rates between the two simulations at $z>1$, with the $w_{0}w_{a}\mathrm{CDM}$ cosmology exhibiting a slower star formation rate due to its lower baryon fraction. Had the Discovery simulations been run with the exact same baryon fraction, we would likely see no substantial offset in star formation rates at high redshift, and discrepancies at lower redshift would become more apparent.
@@ Acknowledgements
We would like to thank the anonymous referee for their helpful comments and suggestions, which have improved the quality of this paper. We thank Salman Habib, Eve Kovacs, Patricia Larsen, Matthew Becker, and other members of the Argonne CPAC group, as well as Johannes Lange, for useful discussions related to this work. We also thank Silvio Rizzi for visualizing the Discovery simulations. Work done at Argonne was supported under the DOE contract DE-AC02-06CH11357. We gratefully acknowledge use of the Bebop, Improv, and Swing supercomputers in the Laboratory Computing Resource Center at Argonne National Laboratory. The Discovery simulations were generated using resources of the Argonne Leadership Computing Facility, a U.S. Department of Energy (DOE) Office of Science user facility at Argonne National Laboratory and is based on research supported by the U.S. DOE Office of Science-Advanced Scientific Computing Research Program. This work was partially supported by the OpenUniverse effort, which is funded by NASA under JPL Contract Task 70-711320, ‘Maximizing Science Exploitation of Simulated Cosmological Survey Data Across Surveys’. The project that gave rise to these results received the support of a fellowship from ”la Caixa” Foundation (ID 100010434). The fellowship code is LCF/BQ/PI23/11970028.